# PolSAM: Polarimetric Scattering Mechanism Informed Segment Anything Model

### v1

@@ V-B Comparison With State-of-the-art Methods
Through a comparison of the red-boxed details in the figure, PolSAM demonstrates a strong capability in capturing edge information between different terrain classes in segmentation results. LSNet [55], which incorporates edge constraints to preserve boundary information during training, outperforms several comparative methods that do not utilize such constraints. However, given the inherent limitations of PolSAR data, including speckle noise and indistinct edges in certain terrain classes, its results remain somewhat unsatisfactory. In contrast, our method achieves better results by more effectively leveraging semantic information to guide the segmentation process, leading to clearer edges and improved overall accuracy.
TABLE III: Ablation studies on module design of PolSAM on PhySAR-Seg-1 and PhySAR-Seg-2
@@
V-C Ablation Studies on Module Design
We perform comprehensive module ablation studies on both datasets to assess the efficacy of the modules. As shown in Table III, we first define the baseline model as the original SAM that takes the pseudo-colored SAR image as the input and only train the mask decoder part. The following models, M1 to M5, all take two inputs. Compared to our method PolSAM (w/ MVD), M2, M3, and M1 are designed to validate the effectiveness of the adapter, FFP, and their combined impact, respectively. It can be seen that the adapter in the PolSAM encoder has a significant influence of the model performance, and the proposed FFP can also improve the result to a certain extent. The effectiveness of the proposed progressive SFP can be verified by comparing the performance of M4. The implementation of the SFP module enhances the overall performance of the model, as evidenced by the results. Compared to M4, our method enhances performance by 3.97% on the PhySAR-Seg-1 dataset, while on the PhySAR-Seg-2 dataset, the enhancement is 5.93%.
@@ V-C Ablation Studies on Module Design
To further demonstrate the validity of our proposed method, we conducted an ablation study on the MVD. In the PolSAM (w/o MVD) experiment, we replaced the MVD with pseudo-colored images. Despite this modification, PolSAM (w/o MVD) still shows significant improvements over the baseline, achieving mIoU increases of 15.35% on the PhySAR-Seg-1 dataset and 8.64% on the PhySAR-Seg-2 dataset, illustrating the soundness of our overall model design. However, when compared to PolSAM (w/ MVD), PolSAM (w/o MVD) falls short. PolSAM (w/ MVD) achieves even greater mIoU increases of 17.06% on PhySAR-Seg-1 and 9.89% on PhySAR-Seg-2 compared to the baseline, highlighting that the MVD, representing the scattering mechanism, provides valuable multidimensional information that enhances segmentation performance, particularly in capturing details.
@@
V-D Discussions on Semantic prompt
In our experiments, the number of sparse prompt $C$ is set to the number of semantic class. Consequently, we visualize $C$ class-specific fusion prompt maps denoted as $\mathbf{V}_{\mathrm{S,D}}^{i},i=1,...,C$ in Fig. 13(b) for discussion. It can be observed that the class-specific fusion prompt map shows significant correlation with each class. The results illustrate the explicit semantic meanings of the learned fusion prompt, which play an important role to guide the decoder to learn the segmentation result.
TABLE IV: Ablation study on the number of sparse prompts (SP)
@@ V-D Discussions on Semantic prompt
Sparse Prompt Parameter Selection. Note that the number of sparse prompts $N$ reflects the degree of semantic correlation. Consequently, we conducted ablation experiments to analyze its impact on model performance. As shown in Table IV, we explored six different values of $N$, each being a multiple of the number of categories in the dataset. This experiment was performed on the PhySAR-Seg-1 dataset, where the model achieved the best performance when $N=6$, which matches the number of classes. However, when $N$ was increased significantly to 10 or 20 times the number of categories, the network’s parameter count grew, but the performance metrics did not consistently improve. This indicates that an excessively large $N$ leads to parameter redundancy, while a value of $N$ smaller than the number of classes can cause semantic confusion. Therefore, we infer that the sparse prompt is closely linked to semantic information, enabling the semantic-level fusion prompt to align semantically during training, which ultimately enhances segmentation results.
TABLE V: Ablation experiment on physical information input forms in the PhySAR-Seg-2 dataset
Metrics Inputs: H/A/Alpha($\alpha$)[56] T[57]
(%) None H/A/$\alpha$ T(6) T(9) H/A/$\alpha$_T MVD
@@
V-E Discussions on Effectiveness of MVD.
Utilizing MVD offers two notable benefits. Initially, the complex information is transformed into a compact format through a series of processing stages. This process allows the data to visually represent the scattering mechanism, effectively conveying the physical information in a way that is intuitively understandable for humans, while also being highly correlated with semantic information. Additionally, MVD requires only a small amount of memory, significantly reducing the storage needed for complex data and improving efficiency in data loading and preprocessing for neural networks.
@@
To validate the effectiveness of MVD, we conducted experiments comparing different polarimetric decomposition features as substitutes for MVD in the PolSAM. The results, shown in Table V, evaluate the performance of four feature combinations, primarily based on the H/A/Alpha decomposition method [56] and the polarization coherence matrix ($T$). We compared the performance metrics, frame per second (FPS), and data memory usage against these alternative decomposition methods.
The H/A/Alpha decomposition method [56] is used to analyse scattering mechanisms in PolSAR data. The entropy (H) and alpha angle ($\alpha$) are calculated using the eigenvalues and eigenvectors of the coherency matrix. This allows for the characterisation of surface, double-bounce, and volume. The entropy (H) is calculated using the formula [$H=-\sum_{i=1}^{3}p_{i}\log_{3}(p_{i})$], where $p_{i}$ represents the probabilities associated with the eigenvalues. The alpha angle ($\alpha$) is calculated using the formula [$\alpha=\sum_{i=1}^{3}p_{i}\alpha_{i}$], where $\alpha_{i}$ represents the corresponding scattering mechanism angles. This technique offers the benefit of offering a comprehensive explanation of scattering mechanisms, but it may be restricted by its susceptibility to noise and the assumption of simplistic scattering models, which may not accurately represent more complex ground features.
The polarization coherence matrix $T$ is derived from the Pauli scattering vector under the reciprocity condition and represents the fully polarimetric information in PolSAR imaging [57]. The matrix is given by:
@@
The real parts of these upper triangular elements form the six-channel T(6) dataset, as shown in the third column of Table V. The full set of real and imaginary components together constitute the nine-channel T(9) dataset, represented in the fourth column. Additionally, the H/A/Alpha_T(12) approach in the fifth column combines the three channels from the H/A/Alpha decomposition with the nine-channel data from the $T$ decomposition, resulting in a comprehensive twelve-channel dataset.
@@
The comparison results in the table first demonstrate the effectiveness of the PolSAM architecture, as all four methods incorporating different polarimetric information show improvements over the baseline, confirming the validity of the model. Furthermore, when comparing these methods to our approach using MVD, the advantages of incorporating MVD become even more evident. Our MVD approach, compared to the second-best method, H/A/Alpha decomposition, shows clear improvements in evaluation metrics, with increases of 3.66% in mAcc, 4.06% in mean F1 score, and 4.3% in mIoU. These results clearly demonstrate that integrating MVD further enhances segmentation performance, highlighting the added value it brings to the PolSAM architecture.
@@
To further evaluate the effectiveness of our method, we conducted a detailed analysis of IoU metrics for each class in the PhySAR-Seg-2 dataset. The results, presented in Table V and visualized in Fig. 14, show that our method using MVD achieves the best results across most classes, particularly for the urban class, which is characterized by double-bounce scattering and has the lowest overall proportion in the dataset.
@@
For loading a batch of 668 image pairs, our method reduces the loading time by 16.52 seconds, 57.11 seconds, and 82.07 seconds compared to the T(6), T(9), and H/A/Alpha_T(12) approaches, respectively. In terms of data storage, MVD occupies only 49 MB of memory, representing just 0.755%, 0.377%, 0.252%, and 0.189% of the capacity utilized by the other four methods. These results highlight the clear advantages of MVD, making it a highly efficient and effective option for polarimetric data processing.
VI Conclusion
@@
In this study, we proposed PolSAM, a novel segmentation framework designed for PolSAR image analysis. PolSAM utilizes the Microwave Vision Data (MVD) representation, which encodes scattering characteristics into a lightweight and interpretable format, offering both visual and physical insights. By integrating features from pseudo-colored SAR images and MVD, PolSAM effectively combines complementary information to guide the segmentation process. The framework incorporates a FFP module for the initial integration of features, supported by an adapter to align representation spaces and address feature differences. Additionally, the SFP module refines segmentation through multi-level feature interactions, enabling the effective incorporation of semantic and physical scattering characteristics. Together, these components enable PolSAM to achieve efficient and accurate segmentation while maintaining high interpretability.
@@
In the future, we aim to explore more robust methods to enhance the adaptability of large models to SAR data for improved utilization efficiency. Additionally, considering the rich physical scattering characteristics inherent in PolSAR data, we will investigate approaches to further enhance segmentation results for PolSAR images.

### v2

@@ 5.2 Comparison With State-of-the-art Methods
As shown in the red-boxed details, PolSAM excels at capturing edge information between different terrain classes. While LSNet Zhou et al. (2023), which incorporates edge constraints, outperforms methods without such constraints, the inherent limitations of PolSAR data, including speckle noise and unclear edges, still affect its performance. In contrast, PolSAM leverages semantic information more effectively, achieving clearer edges and superior overall segmentation accuracy.
Table 3: Ablation studies on module design of PolSAM on PhySAR-Seg-1 and PhySAR-Seg-2.
@@
5.3 Ablation Studies on Module Design
We perform comprehensive module ablation studies on both datasets to assess module efficacy. As shown in Table 3, we define the baseline model as original SAM using pseudo-colored SAR images as input while training only the mask decoder. Models M1 to M5 all utilize two inputs. Compared to our method PolSAM (w/ MVD), M2, M3, and M1 validate the effectiveness of the adapter, FFP, and their combined impact respectively. The adapter in the PolSAM encoder significantly influences model performance, while the proposed FFP also improves results considerably. The effectiveness of the proposed progressive SFP is verified by comparing M4 performance. SFP module implementation enhances overall model performance, as evidenced by the results. Compared to M4, our method enhances performance by 3.97% on PhySAR-Seg-1 and 5.93% on PhySAR-Seg-2.
@@ 5.3 Ablation Studies on Module Design
To further demonstrate our proposed method’s validity, we conducted an ablation study on MVD. In the PolSAM (w/o MVD) experiment, we replaced MVD with pseudo-colored images. Despite this modification, PolSAM (w/o MVD) still shows significant improvements over the baseline, achieving mIoU increases of 15.35% on PhySAR-Seg-1 and 8.64% on PhySAR-Seg-2, illustrating the soundness of our overall model design. However, PolSAM (w/o MVD) falls short compared to PolSAM (w/ MVD). The latter achieves greater mIoU increases of 17.06% on PhySAR-Seg-1 and 9.89% on PhySAR-Seg-2 compared to the baseline, highlighting that MVD, representing scattering mechanisms, provides valuable multidimensional information that enhances segmentation performance, particularly in detail capture.
5.4 Discussions on Semantic prompt
@@
In our experiments, the number of sparse prompts $C$ is set to the number of semantic classes. Consequently, we visualize $C$ class-specific fusion prompt maps denoted as $\mathbf{V}_{\text{S,D}}^{i},i=1,...,C$ in Fig. 12 (b) for discussion. It can be observed that class-specific fusion prompt maps show significant correlation with their respective classes. The results illustrate the explicit semantic meanings of learned fusion prompts, which play important roles in guiding the decoder toward optimal segmentation results.
Sparse Prompt Parameter Selection. The number of sparse prompts $N$ reflects the degree of semantic correlation. To evaluate its impact on model performance, we conducted ablation experiments using six different $N$ values, each being a multiple of the dataset’s category number. As shown in Table 4, optimal performance on PhySAR-Seg-1 was achieved when $N=6$, matching the class number. Increasing $N$ to 10 or 20 times the category number increased the model’s parameter count without yielding consistent performance improvements, suggesting that excessively large $N$ leads to parameter redundancy, while $N$ values smaller than the class number cause semantic confusion. These results indicate that sparse prompts, when aligned with semantic information, facilitate effective semantic-level fusion, improving segmentation accuracy.
Table 4: Ablation study on the number of sparse prompts (SP).
@@
5.5 Discussions on Effectiveness of MVD.
Utilizing MVD offers two notable benefits. First, complex information is transformed into compact format through a series of processing stages. This process enables data to visually represent scattering mechanisms, effectively conveying physical information in an intuitively understandable manner while maintaining high correlation with semantic information. Second, MVD requires minimal memory, significantly reducing storage requirements for complex data and improving efficiency in data loading and preprocessing for neural networks.
To validate MVD effectiveness, we conducted experiments comparing different polarimetric decomposition features as MVD substitutes in PolSAM. The results in Table 5 evaluate four feature combinations, primarily based on H/A/Alpha decomposition Cloude and Pottier (1997) and the polarization coherence matrix ($T$). We compared performance metrics, frames per second (FPS), and data memory usage against these alternative decomposition methods.
The H/A/Alpha decomposition method Cloude and Pottier (1997) is used to analyse scattering mechanisms in PolSAR data. The entropy (H) and alpha angle ($\alpha$) are calculated using the eigenvalues and eigenvectors of the coherency matrix. This allows for the characterisation of surface, double-bounce, and volume. The entropy (H) is calculated using the formula $H=-\sum_{i=1}^{3}p_{i}\log_{3}(p_{i})$], where $p_{i}$ represents the probabilities associated with the eigenvalues. The alpha angle ($\alpha$) is calculated using the formula $\alpha=\sum_{i=1}^{3}p_{i}\alpha_{i}$], where $\alpha_{i}$ represents the corresponding scattering mechanism angles. This technique offers the benefit of offering a comprehensive explanation of scattering mechanisms, but it may be restricted by its susceptibility to noise and the assumption of simplistic scattering models, which may not accurately represent more complex ground features.
Table 5: Ablation experiment on physical information input forms in the PhySAR-Seg-2 dataset.
Metrics Inputs: H/A/Alpha($\alpha$) / T
[…]
@@
The polarization coherence matrix $T$ is derived from the Pauli scattering vector under the reciprocity condition and represents the fully polarimetric information in PolSAR imaging Chen and Tao (2018). The matrix is given by:
@@
The real parts of these upper triangular elements form the six-channel T(6) dataset, as shown in the third column of Table 5. The full set of real and imaginary components together constitute the nine-channel T(9) dataset, represented in the fourth column. Additionally, the H/A/Alpha_T(12) approach combines the three channels from the H/A/Alpha decomposition with the nine-channel data from the $T$ decomposition, resulting in a comprehensive twelve-channel dataset.
@@
The comparison results first demonstrate PolSAM architecture effectiveness, as all four methods incorporating different polarimetric information show improvements over the baseline, confirming model validity. Furthermore, when comparing these methods to our MVD approach, the advantages of incorporating MVD become more evident. Our MVD approach, compared to the second-best method (H/A/Alpha decomposition), shows clear improvements in evaluation metrics, with increases of 3.66% in mAcc, 4.06% in mean F1 score, and 4.3% in mIoU. These results clearly demonstrate that integrating MVD further enhances segmentation performance, highlighting its added value to the PolSAM architecture.
@@
To further evaluate our method’s effectiveness, we conducted detailed analysis of IoU metrics for each class in the PhySAR-Seg-2 dataset. The results, presented in Table 5 and visualized in Fig. 13, show that our MVD-based method achieves optimal performance across most classes, particularly the urban class, characterized by double-bounce scattering and having the lowest dataset proportion.
@@
For loading a batch of 668 image pairs, our method reduces loading time by 16.52 seconds, 57.11 seconds, and 82.07 seconds compared to T(6), T(9), and H/A/Alpha_T(12) approaches respectively. Regarding data storage, MVD occupies only 49 MB of memory, representing merely 0.755%, 0.377%, 0.252%, and 0.189% of the capacity utilized by the other four methods. These results highlight MVD’s clear advantages, making it a highly efficient and effective option for polarimetric data processing.
6 Conclusion
@@
In this study, we propose PolSAM, a novel segmentation framework for PolSAR image analysis. PolSAM utilizes Microwave Vision Data (MVD) representation, which encodes scattering characteristics into a lightweight and interpretable format, offering both visual and physical insights. By integrating features from pseudo-colored SAR images and MVD, PolSAM effectively combines complementary information to guide the segmentation process. The framework incorporates an FFP module for initial feature integration, supported by adapters to align representation spaces and address feature differences. Additionally, the SFP module refines segmentation through multi-level feature interactions, enabling effective incorporation of semantic and physical scattering characteristics. Together, these components enable PolSAM to achieve efficient and accurate segmentation while maintaining high interpretability.
@@
Future work will explore more robust methods to enhance large model adaptability to SAR data for improved utilization efficiency. Additionally, considering the rich physical scattering characteristics inherent in PolSAR data, we will investigate approaches to further enhance PolSAR image segmentation results.
7 Acknowledgment
This work was supported by the National Natural Science Foundation of China (Grant No. 62571443); the National Key Laboratory of Microwave Imaging Technology; the Natural Science Basic Research Program of Shaanxi (Program No. 2025JC-QYXQ-032); and the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2025A1515011368).